### 05 Mar 2026 v1

Abstract
We present guidelines for deriving new Nitsche Finite Element Methods to enforce equality and inequality constraints that act on the value of the unknown mechanical quality. We first formulate the problem as a stabilized finite element method for the saddle point formulation where a Lagrange multiplier enforces the underlying constraint. The Nitsche method is then presented in a general minimization form, suitable for nonlinear finite element methods and allowing straightforward computational implementation with automatic differentation. To validate these ideas, we present Nitsche formulations for a range of problems in solid mechanics and give numerical evidence of the convergence rates of the Nitsche method.
keywords:
Nitsche method , finite element method , inequality constraints , contact problems , stabilization
MSC:
65N30 , 74M15 , 49J40
journal: International Journal of Engineering Science11footnotetext: This work was supported by the Research Council of Finland (Flagship of Advanced Mathematics for Sensing Imaging and Modelling grant 359181) and the Portuguese government through FCT (Fundação para a Ciência e a Tecnologia), I.P., under the project UIDB/04459/2025.
[…]
[mech]organization=Department of Mechanical Engineering, Aalto University, addressline=P.O. Box 11100, postcode=00076 Aalto, city=Espoo, country=Finland
[…]
[msa]organization=Department of Mathematics and Systems Analysis, Aalto University, city=Espoo, country=Finland
[…]
[ist]organization=CAMGSD/Departamento de Matemática, Instituto Superior Técnico, Universidade de Lisboa, addressline=Av. Rovisco Pais 1, postcode=1049-001, city=Lisbon, country=Portugal
{highlights}
We derive a novel interpretation of the Nitsche Finite Element Method to solve constrained problems in solid mechanics
We use the new formulation to solve several contact problems and validate the convergence rates numerically
The computational implementation utilizes a general minimization form and automatic differentation handles the variational derivatives
1 Introduction
In the 1970s, Nitsche [nitsche] introduced a method for enforcing the Dirichlet boundary conditions in the weak formulation of the finite element method. The benefits of the Nitsche method over its alternatives, the penalty method [courant1943variational] and the mixed Lagrange multiplier method [babuvska1973finite], include consistency, symmetry and positive definiteness, stability, and conditioning; cf. [gustafsson2025]. Hence, the ideas of Nitsche have later found applications also in domain decomposition [becker2003finite], discontinuous Galerkin methods [hansbo2002discontinuous], unfitted finite element and fictitious domain methods [burman2015cutfem], Robin boundary conditions [juntunen2009nitsche], Kirchhoff plates [gustafsson2021nitsche], elastic contact problems [chouly2013nitsche], friction problems [chouly2014adaptation], obstacle problems [gustafsson2017mixed], elastoplastic torsion problems [chouly2023nitsche], and plate contact problems [gustafsson2017nitsche], to name a few.
In modern literature, the Nitsche method is often presented as a consistency correction to the penalty method, i.e., to make sure that the exact solution satisfies the discrete weak form, by including additional terms; cf., e.g.,  [benzaken2024constructing, gustafsson2025]. This fails to answer to the obvious follow-up question: how to generalize the method to other problems and constraints? Consequently, we take a different viewpoint based on the saddle point formulation and present guidelines to derive optimal Nitsche methods for both equality and inequality constraints. These guiding principles apply to arbitrary constrained minimization problems, where the constraint acts on the values of the unknown. The resulting energy minimization form is particularly well-suited for solving nonlinear problems numerically. We follow the guidelines in practice by deriving and implementing novel methods for problems such as inequality boundary condition for plates, membrane-solid contact problem, and plate-plate contact problem.
Our formulation is based on a reinterpretation of the Nitsche method by Stenberg [stenberg1995]. More precisely, on the idea that the Nitsche method corresponds to a conforming stabilized Lagrange multiplier method (in the sense of residual Barbosa–Hughes stabilization [barbosa1992circumventing]). For such stabilized methods, an element-wise elimination of the Lagrange multiplier will lead to the equivalent Nitsche method; see also [chouly2013nitsche, burman2017galerkin, chouly2023finite, burman2023augmented]. This connection was exploited in the error analysis in Gustafsson [gustafsson2018finite] and consequently it became evident that the stabilized formulation leads to a family of methods with constraints of rather general form.
Let us summarize the prior theoretical findings for existing Nitsche methods: (1) stability becomes automatic if correct residuals of the Lagrange multiplier are included in the formulation; (2) optimal convergence rates are obtained if the scaling of the residual with respect to the mesh and the material parameters correctly mimics the continuous Sobolev norm of the Lagrange multiplier.
In this paper, built on these observations, we explain how to define and use general Nitsche methods, present examples of their adaptation to novel physical problems, and show numerical evidence of their convergence. A more careful numerical analysis, i.e., proofs of stability and error estimates, is beyond the scope of this work. The remainder of the paper is organized as follows. In Section 2, we review the original Nitsche method for the Dirichlet boundary condition $u=g$. In Section 3, we demonstrate how the method is modified to satisfy the inequality constraint $u\geq g$. In Section 4, we describe a general form of the Nitsche method and how it can be used to derive new methods. In Section 5, we present some methods that exist in the literature and are of this general form. In Section 6, we derive novel Nitsche methods from the general form and show numerical evidence of their performance.
2 Background: Dirichlet boundary conditions
For better context, we briefly recall the original method of Nitsche [nitsche]. Consider the model problem
$-\nabla\cdot\kappa\nabla u$ $=f,\quad\text{in $\Omega$},$ (1)
$u$ $=g,\quad\text{on $\partial\Omega$},$
where $f\in L^{2}(\Omega)$ is the problem-specific load, $g\in H^{1/2}(\partial\Omega)$ is the Dirichlet boundary condition and $\kappa>0$ is a given parameter. Note that Nitsche [nitsche] considered only the case $\kappa=1$, i.e., normalized units.
Problem (1) is solved in a finite element space $V_{h}\subset H^{1}(\Omega)$ where $h$ is the mesh parameter. The classical Nitsche method for the Dirichlet boundary condition $u=g$ reads: find $u_{h}\in V_{h}$ such that
$\int_{\Omega}\kappa\nabla u_{h}\cdot\nabla v_{h}+\int_{\partial\Omega}\frac{1}{\gamma}u_{h}v_{h}-\int_{\partial\Omega}\kappa\frac{\partial u_{h}}{\partial n}v_{h}-\int_{\partial\Omega}u_{h}\,\kappa\frac{\partial v_{h}}{\partial n}$ (2)
$\quad=\int_{\Omega}fv_{h}+\int_{\partial\Omega}\frac{1}{\gamma}gv_{h}-\int_{\partial\Omega}g\,\kappa\frac{\partial v_{h}}{\partial n}\qquad\forall v_{h}\in V_{h}.$
Here $\gamma$ is an $L^{2}(\Omega)$ function whose restriction to each element $K$ is the product of a dimensionless stabilization parameter $\alpha>0$ and the local mesh parameter $h_{K}$, divided by $\kappa$, i.e.
$\gamma|_{K}=\frac{\alpha h_{K}}{\kappa}.$
According to Stenberg [stenberg1995], the Nitsche method corresponds to using stabilized finite element method to discretize the saddle point problem: find $(u,\lambda)\in H^{1}(\Omega)\times H^{-1/2}(\partial\Omega)$ such that
$\mathcal{L}(u,\lambda)=\inf_{v\in H^{1}(\Omega)}\sup_{\mu\in H^{-1/2}(\partial\Omega)}\mathcal{L}(v,\mu).$ (3)
The Lagrangian in (3) is defined as
$\mathcal{L}(u,\lambda)=J(u)-\langle u-g,\lambda\rangle,$ (4)
where
$J(u)=\frac{1}{2}\int_{\Omega}\kappa\nabla u\cdot\nabla u-\int_{\Omega}fu$ (5)
is the energy functional of the original problem (1) [brezis2011functional], $\langle\cdot,\cdot\rangle$ denotes the duality pairing between $H^{1/2}(\partial\Omega)$ and $H^{-1/2}(\partial\Omega)$, and $\lambda$ is a Lagrange multiplier. We are able to derive the strong formulation of problem (3) to reveal that $\lambda=\kappa\frac{\partial u}{\partial n}$.
The Nitsche method (2) corresponds to the stabilized discrete saddle point problem: find $(u_{h},\lambda_{h})\in V_{h}\times Q_{h}$ such that
$\mathcal{L}_{h}(u_{h},\lambda_{h})=\inf_{v_{h}\in V_{h}}\sup_{\mu_{h}\in Q_{h}}\mathcal{L}_{h}(v_{h},\mu_{h})$ (6)
where the stabilized Lagrangian $\mathcal{L}_{h}$ is defined as
$\mathcal{L}_{h}(v_{h},\mu_{h})=\Big\{\mathcal{L}(v_{h},\mu_{h})-\int_{\partial\Omega}\frac{\gamma}{2}\Big(\mu_{h}-\kappa\frac{\partial v_{h}}{\partial n}\Big)^{2}\Big\}.$ (7)
The space $Q_{h}\subset H^{-1/2}(\partial\Omega)$ denotes a set of elementwise polynomials that are globally discontinuous. The variational formulation of problem (6) reads: find $(u_{h},\lambda_{h})\in V_{h}\times Q_{h}$ such that
$\int_{\Omega}\kappa\nabla u_{h}\cdot\nabla v_{h}-\int_{\partial\Omega}\lambda_{h}v_{h}-\int_{\partial\Omega}u_{h}\mu_{h}$ (8)
$\quad-\int_{\partial\Omega}\gamma\Big(\lambda_{h}-\kappa\frac{\partial u_{h}}{\partial n}\Big)\Big(\mu_{h}-\kappa\frac{\partial v_{h}}{\partial n}\Big)=\int_{\Omega}fv_{h}-\int_{\partial\Omega}g\mu_{h}$
for every $(v_{h},\mu_{h})\in V_{h}\times Q_{h}$. This method is stable for any pair of finite element spaces $V_{h}$ and $Q_{h}$ if $0<\alpha<C_{I}$ where $C_{I}$ is a constant of an inverse estimate, cf. [stenberg1995].
Following the steps laid out in Stenberg [stenberg1995], we can show that an elimination of the Lagrange multiplier element-by-element leads to the formulation (2). Assuming that $\kappa$ is an element-wise constant and choosing $v_{h}=0$ leads to
$\int_{\partial\Omega}u_{h}\mu_{h}-\int_{\partial\Omega}\gamma\Big(\lambda_{h}-\kappa\frac{\partial u_{h}}{\partial n}\Big)\mu_{h}=-\int_{\partial\Omega}g\mu_{h}\quad\forall\mu_{h}\in Q_{h}.$ (9)
Denoting by $\pi_{h}$ the $L^{2}$ projection onto $Q_{h}$, this reads
$\pi_{h}u_{h}-\gamma\Big(\lambda_{h}-\kappa\pi_{h}\frac{\partial u_{h}}{\partial n}\Big)=-\pi_{h}g$ (10)
which gives
$\lambda_{h}=\kappa\pi_{h}\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(\pi_{h}u_{h}-\pi_{h}g).$ (11)
Let $V_{h}$ correspond to the space of continuous piecewise polynomial functions of degree $p\geq 1$. If $Q_{h}$ is chosen as discontinuous piecewise polynomials of degree $p$ or greater, then $\pi_{h}$ reduces to the identity mapping, i.e.
$\lambda_{h}=\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g).$ (12)
Now we can eliminate the Lagrange multiplier from the discrete saddle point problem. After substituting $\lambda_{h}$ to the stabilized Lagrangian $\mathcal{L}_{h}$ we obtain
$J(u_{h})$ $-\int_{\partial\Omega}\Big(\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)\Big)(u_{h}-g)$
$\quad-\int_{\partial\Omega}\frac{\gamma}{2}\Big(\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)-\kappa\frac{\partial u_{h}}{\partial n}\Big)^{2}$
which simplifies to
$J(u_{h})$ $-\int_{\partial\Omega}\kappa\frac{\partial u_{h}}{\partial n}(u_{h}-g)+\int_{\partial\Omega}\frac{1}{\gamma}(u_{h}-g)^{2}$
$\quad-\int_{\partial\Omega}\frac{1}{2\gamma}(u_{h}-g)^{2}$
and finally to
$J_{h}(u_{h})=J(u_{h})-\int_{\partial\Omega}\kappa\frac{\partial u_{h}}{\partial n}(u_{h}-g)+\frac{1}{2}\int_{\partial\Omega}\frac{1}{\gamma}(u_{h}-g)^{2}.$ (13)
The classical Nitsche method (2) corresponds to the following minimization problem [nitsche]: find $u_{h}\in V_{h}$ such that
$J_{h}(u_{h})=\inf_{v_{h}\in V_{h}}J_{h}(v_{h}).$ (14)
Note that by adding and subtracting $\int_{\partial\Omega}\frac{\gamma}{2}(\kappa\frac{\partial u_{h}}{\partial n})^{2}$ to complete the square we can rewrite $J_{h}$ in (13) as
$J(u_{h})+\int_{\partial\Omega}\frac{\gamma}{2}\Big(\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)\Big)^{2}-\int_{\partial\Omega}\frac{\gamma}{2}\Big(\kappa\frac{\partial u_{h}}{\partial n}\Big)^{2}.$
This form can be compared to the penalty method which simply minimizes the penalized functional:
$J(u_{h})+\frac{1}{2\varepsilon}\int_{\partial\Omega}(u_{h}-g)^{2}.$
The penalty method approximates the Robin boundary condition
$\kappa\frac{\partial u}{\partial n}=\frac{1}{\varepsilon}(u-g),$ (15)
which satisfies $u-g\rightarrow 0$ when $\varepsilon\rightarrow 0$. Unfortunately, small enough values of $\varepsilon$ lead to ill-conditioned linear systems [gustafsson2025].
3 Inequality constraints: Signorini’s problem
For inequality constraints, we first consider problem (1) with the inequality constraint $u\geq g$ on the boundary $\partial\Omega$. The resulting problem
$-\nabla\cdot\kappa\nabla u$ $=f,\quad\text{in $\Omega$},$ (16)
$u\geq g,~\kappa\frac{\partial u}{\partial n}\geq 0,~(u-g)\kappa\frac{\partial u}{\partial n}$ $=0,\quad\text{on $\partial\Omega$},$
is often referred to as the scalar Signorini’s or Poisson–Signorini problem [gustafsson2019nitsche]. The corresponding continuous saddle point formulation of (16) is: find $(u,\lambda)\in H^{1}(\Omega)\times H^{-1/2}(\partial\Omega)$, $\lambda\geq 0$, such that
$\mathcal{L}(u,\lambda)=\inf_{v\in H^{1}(\Omega)}\sup_{\mu\in H^{-1/2}(\partial\Omega),\,\mu\geq 0}\mathcal{L}(v,\mu),$ (17)
where the Lagrangian $\mathcal{L}$ was defined in (4). The connection between the two formulations is again given by $\lambda=\kappa\frac{\partial u}{\partial n}$. After residual stabilization, the finite element formulation reads: find $(u_{h},\lambda_{h})\in V_{h}\times Q_{h}$, $\lambda_{h}\geq 0$, such that
$\mathcal{L}_{h}(u_{h},\lambda_{h})=\inf_{v_{h}\in V_{h}}\sup_{\mu_{h}\in Q_{h},\,\mu_{h}\geq 0}\mathcal{L}_{h}(v_{h},\mu_{h}),$ (18)
where the stabilized Lagrangian $\mathcal{L}_{h}$ was defined in (7). The error analysis of formulation (18) was given in [gustafsson2019nitsche] where it is concluded, among other things, that the formulation is stable for any conforming $Q_{h}$.
The corresponding variational inequality can be written as: find $\lambda_{h}\in Q_{h}$, $\lambda_{h}\geq 0$, so that
$\int_{\Omega}\kappa\nabla u_{h}\cdot\nabla v_{h}-\int_{\partial\Omega}\lambda_{h}v_{h}+\int_{\partial\Omega}\gamma\Big(\lambda_{h}-\kappa\frac{\partial u_{h}}{\partial n}\Big)\,\kappa\frac{\partial v_{h}}{\partial n}$ $=\int_{\Omega}fv_{h},$
$\int_{\partial\Omega}\Big(\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)-\lambda_{h}\Big)(\mu_{h}-\lambda_{h})$ $\leq 0,$
for every $(v_{h},\mu_{h})\in V_{h}\times Q_{h}$, $\mu_{h}\geq 0$. The second inequality implies that $\lambda_{h}$ is the orthogonal $L^{2}$ projection of $\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)$ onto the set of nonnegative functions in $Q_{h}$ [brezis2011functional, Theorem 5.2]. Since $Q_{h}$ is arbitrary, we may choose $Q_{h}=L^{2}(\partial\Omega)$ and then the orthogonal projection is given explicitly by the maximum operator, i.e.,
$\lambda_{h}=\max(\kappa\tfrac{\partial u_{h}}{\partial n}-\tfrac{1}{\gamma}(u_{h}-g),0)=:(\kappa\tfrac{\partial u_{h}}{\partial n}-\tfrac{1}{\gamma}(u_{h}-g))_{+}.$ (19)
In order to obtain the Nitsche method, we again substitute the discrete Lagrange multiplier $\lambda_{h}$ from (19) to (18) and rearrange the terms to obtain: find $u_{h}\in V_{h}$ which minimizes
$J(u_{h})+\int_{\partial\Omega}\frac{\gamma}{2}\Big(\kappa\frac{\partial u_{h}}{\partial n}-\frac{1}{\gamma}(u_{h}-g)\Big)_{+}^{2}-\int_{\partial\Omega}\frac{\gamma}{2}\Big(\kappa\frac{\partial u_{h}}{\partial n}\Big)^{2}.$ (20)
Note that the corresponding method for the equality constraint can be obtained by dropping the maximum operator (subscript ++).
4 Structure of the method
We now generalize the method to arbitrary inequality constraints of the form $\beta(u_{h})\geq 0$. The method (20) has the following main ingredients: the energy of the source problem, i.e., $J$; the constraint to impose, i.e., $\beta(u_{h})=u_{h}-g\geq 0$; the corresponding definition of the continuous Lagrange multiplier, i.e., $\lambda(u_{h})=\kappa\frac{\partial u_{h}}{\partial n}$; and the correct scaling of the stabilization parameter $\gamma$ with respect to the mesh parameter $h$ and any material parameters, i.e., $\gamma(h,\kappa)|_{K}=\tfrac{\alpha h_{K}}{\kappa}$. In conclusion, the general form of the Nitsche method can be written as follows: find $u_{h}\in V_{h}$ that minimizes the functional
$J(u_{h})+\int_{\Gamma}\frac{\gamma(h,\kappa)}{2}\Big(\lambda(u_{h})-\frac{1}{\gamma(h,\kappa)}\beta(u_{h})\Big)_{+}^{2}-\int_{\Gamma}\frac{\gamma(h,\kappa)}{2}\lambda(u_{h})^{2}.$ (21)
The final method is formulated as a minimization problem instead of a variational one as it is the starting point for our implementation of the method; see Section 6 and Appendix A. In particular, the calculation of the second functional derivative of (21) is performed automatically in our computational implementation of the Nitsche method.
Remark 1
We emphasize that the general form (21) was obtained following the derivation of (20). Equivalently, it can be deduced starting from a similar minimization form existing in the literature for other physical problems; see, e.g., [chouly2013nitsche, p. 135] for elastic contact problems and [burman2017galerkin, p. 364] for the obstacle problem. The form (21) is also present in [burman2023augmented] where the authors discuss an alternative approach using augmented Langrangian methods.
Hence, for new problems written in the general form (21), with energy $J(u_{h})$ and finite element space $V_{h}$, it is necessary to define:
  1. 1.
    A constraint $\beta=\beta(u_{h})\geq 0$ and the subset $\Gamma\subset\Omega$ where the constraint can become active.
  2. 2.
    An expression for the continuous Lagrange multiplier (i.e., the contact force) as $\lambda=\lambda(u_{h})$.
  3. 3.
    An expression for the parameter $\gamma=\gamma(h,\kappa)$.
In order to find an expression for $\lambda=\lambda(u_{h})$, it is necessary to derive the strong formulation of the continuous saddle point problem. This is usually obtained by first deriving the variational formulation and then integrating by parts to find the corresponding differential equations and boundary conditions, some of which will depend directly on $\lambda$. For contact-type constraints between two domains, prior research suggests that the parameter $\gamma$ and, hence, the expression for $\lambda$ should be defined on the less stiff side; cf., e.g., [gustafsson2019error].
Physically, equation (21) shows that an important role of $\gamma$ is to scale the units of displacement in $\beta$ to the units of force in $\lambda$. More precisely, $\lambda$ should be proportional to $\frac{\beta}{\gamma}$ with the proportionality constant independent of $h$ and $\kappa$. Therefore, the material parameters of the problem must be included in $\gamma$. Mathematically speaking, a correct scaling of $\gamma$ is necessary to prove the uniform stability of the discrete saddle point formulation using an inverse inequality; cf., e.g., [gustafsson2017mixed, gustafsson2017nitsche, gustafsson2019error].
Finally, we wish to remark that the penalty method can be recovered from this formulation by removing the Lagrange multiplier terms (formally, $\lambda=0$), and that the Nitsche method for equality constraints is obtained simply by dropping the maximum operator (subscript ++). The solution of the minimization problem (21) using Newton’s method is explained in Appendix A.
5 Some existing methods
5.1 Membrane obstacle problem
Let $\Omega$ be a polygonal/polyhedral domain representing a Poisson membrane in its undeformed state. Minimization of the energy
$J(u)=\frac{1}{2}\int_{\Omega}\kappa\nabla u\cdot\nabla u-\int_{\Omega}fu$ (22)
corresponds to solving the Poisson equation $\nabla\cdot(\kappa\nabla u)=f$ with $u\in H^{1}_{0}(\Omega)$. Let now the deflection of the membrane $u$ be constrained by the rigid obstacle $g$ below the membrane, i.e., $u\geq g$. The Lagrange multiplier corresponding to the constraint $u-g\geq 0$ (in $\Omega$) is given by $\lambda=-\nabla\cdot(\kappa\nabla u)-f=:-Lu-f$, cf. [gustafsson2017mixed, burman2017galerkin]. The scaling is given by $\gamma|_{K}=\frac{\alpha h_{K}^{2}}{\kappa}$. Hence, the Nitsche functional reads
$J(u_{h})+\int_{\Omega}\frac{\gamma}{2}\left(-L_{h}u_{h}-f-\frac{1}{\gamma}(u_{h}-g)\right)_{+}^{2}-\int_{\Omega}\frac{\gamma}{2}\left(-L_{h}u_{h}-f\right)^{2},$ (23)
where $L_{h}$ denotes an element-wise evaluation of $L$.
5.2 Two-body linearized elastic contact
Let $\Omega_{1}$ and $\Omega_{2}$ be two polygonal/polyhedral domains that are initially at contact on $\Gamma=\partial\Omega_{1}\cap\partial\Omega_{2}$. The energy of the two-body linear elastic source problem reads
$J((\bm{w}_{1},\bm{w}_{2}))=\sum_{i=1}^{2}\left\{\frac{1}{2}\int_{\Omega_{i}}\bm{\sigma}(\bm{w}_{i}):\bm{\varepsilon}(\bm{w}_{i})-\int_{\Omega_{i}}\bm{f}\cdot\bm{w}_{i}\right\},$ (24)
where
$\bm{\sigma}(\bm{w})$ $=2\mu_{L}\,\bm{\varepsilon}(\bm{w})+\lambda_{L}\mathrm{tr}\,\bm{\varepsilon}(\bm{w})\bm{I},$ (25)
$\bm{\varepsilon}(\bm{w})$ $=\frac{1}{2}(\nabla\bm{w}+\nabla\bm{w}^{T}),$ (26)
and $\bm{f}$ is a given body load and $(\mu_{L},\lambda_{L})$ are the Lamé material parameters.
The physical non-penetration constraint on $\Gamma$ reads $\bm{u}_{1}\cdot\bm{n}_{1}+\bm{u}_{2}\cdot\bm{n}_{2}\leq 0$ where $\bm{n}_{i}$ is the outward normal vector on $\partial\Omega_{i}$, $i=1,2$. The Lagrange multiplier corresponding to the constraint $(\bm{u}_{2}-\bm{u}_{1})\cdot\bm{n}_{1}\geq 0$ is given, e.g., by $\lambda=-\bm{\sigma}(\bm{u}_{1})\bm{n}_{1}\cdot\bm{n}_{1}$. Note that the choice to use the normal traction on $\partial\Omega_{1}$ instead of $\partial\Omega_{2}$ is now arbitrary. In general, the stabilization should be on the less stiff body or the body with smaller elements; cf. [gustafsson2019error]. The scaling is given by $\gamma|_{K}=\frac{\alpha h_{K}}{\mu_{L}}$. The Nitsche functional [chouly2015symmetric] reads
$J((\bm{u}_{1,h},\bm{u}_{2,h}))$ $+\int_{\Gamma}\frac{\gamma}{2}\left(-\bm{\sigma}(\bm{u}_{1,h})\bm{n}_{1}\cdot\bm{n}_{1}-\frac{1}{\gamma}(\bm{u}_{2,h}\cdot\bm{n}_{1}-\bm{u}_{1,h}\cdot\bm{n}_{1})\right)_{+}^{2}$ (27)
$\quad-\int_{\Gamma}\frac{\gamma}{2}\left(-\bm{\sigma}(\bm{u}_{1,h})\bm{n}_{1}\cdot\bm{n}_{1}\right)^{2}.$
6 Novel methods
We now present some novel methods based on formulation (20). More precisely, we consider the following problems:
  1. 1.
    Two-membrane contact problem,
  2. 2.
    Membrane against solid contact problem,
  3. 3.
    Plate against plate contact problem,
  4. 4.
    Kirchhoff plate with inequality boundary constraint.
For each problem, we present the Nitsche formulation and numerically investigate its convergence rate. The numerical results are computed using scikit-fem [gustafsson2020scikit] which is used to solve the nonlinear variational problems using Newton’s method. For completeness, we describe how to manipulate the energy minimization problems for Newton’s method in Appendix A. The programming implementation utilizes automatic differentiation with JAX [bradbury2021jax] so that it is not necessary to manually calculate the functional derivatives involved in Newton’s method. The source code is available in [sourcepackage] for reproducing the numerical results.
Convergence rates are estimated via uniform mesh refinement. In the absence of an analytical solution, we estimate the convergence rate by evaluating the error between two subsequent discrete solutions corresponding to mesh parameters $2h$ and $h$. Assuming that the error satisfies $\|u-u_{h}\|\leq Ch^{p}$ in the energy norm for some $C>0$ and a given polynomial degree $p$ of the finite element basis, we can show that the difference between two subsequent discrete solutions should be bounded by the same asymptotic rate:
$\|u_{h}-u_{2h}\|\leq\|u-u_{h}\|+\|u-u_{2h}\|\leq C_{1}h^{p}+C_{2}(2h)^{p}\leq C_{3}h^{p},\quad C_{1},C_{2},C_{3}>0.$
6.1 Two-membrane contact problem
Figure 1: Numerical solution for two membranes in contact using the Nitsche method.
Let $\Omega\subset\mathbb{R}^{2}$ be a polygonal domain describing the shape of two structurally similar membranes modeled by the Poisson equation that are initially $g$ (units of distance) apart. We denote the vertical displacements of the membranes by $u_{1},u_{2}:\Omega\to\mathbb{R}$ and their tensions by $\kappa_{1},\kappa_{2}>0$, respectively. The membranes are initially separated by a gap $g>0$. The external loading $f_{1},f_{2}:\Omega\to\mathbb{R}$ can cause the membranes to come in contact. Therefore, we introduce a contact constraint $u_{1}-u_{2}\leq g$ in $\Omega$ and a corresponding contact pressure $\lambda:\Omega\to\mathbb{R}$ to prevent penetration.
The strong form of the problem reads as: find the displacements $u=(u_{1},u_{2})$ and the contact pressure $\lambda$ between the membranes such that
$-\nabla\cdot(\kappa_{1}\nabla u_{1})+\lambda$ $=f_{1}$ $\qquad\text{in }\Omega,$ (28)
$-\nabla\cdot(\kappa_{2}\nabla u_{2})-\lambda$ $=f_{2}$ $\qquad\text{in }\Omega,$
$u_{1}=0,\quad u_{2}$ $=0$ $\qquad\text{on }\partial\Omega,$
$u_{1}-u_{2}\leq g,\quad\lambda$ $\geq 0,\quad\lambda(u_{2}-u_{1}+g)=0$ $\qquad\text{in }\Omega.$
Remark 2
We define the problem such that the unknowns are the displacements of the membranes. Other authors, e.g. [belgacem2009membranes, ben2012unilateral], solve for the displaced coordinates of the membranes.
Explicitly, the second membrane is on top, initially $g$ units higher than the first membrane. The constraint $u_{1}-u_{2}\leq g$ prevents penetration by stating that the relative displacement of the membranes does not exceed the initial gap $g$. The contact pressure $\lambda$ appears with opposite signs in the equilibrium equations to represent the opposing forces in contact. The displacements $u_{1}$ and $u_{2}$ are zero on the boundary and we shall numerically approximate them in $\Omega$. We analytically solve and substitute the discrete $\lambda$ in the Nitsche method.
To derive the Nitsche method, we first identify the total potential energy of the unconstrained system, i.e. ignoring the contact force $\lambda$. Multiplying the first equilibrium equation with a test function $v_{1}\in H_{0}^{1}(\Omega)$ and integrating by parts yields the weak form
$\int_{\Omega}-\nabla\cdot(\kappa_{1}\nabla u_{1})v_{1}=\int_{\Omega}\kappa_{1}\nabla u_{1}\cdot\nabla v_{1}$ $=\int_{\Omega}f_{1}v_{1},$
$\implies\int_{\Omega}\kappa_{1}\nabla u_{1}\cdot\nabla v_{1}-\int_{\Omega}f_{1}v_{1}$ $=0.$
It is clear that the first term is the variation (Gateaux derivative) of $\frac{1}{2}\int_{\Omega}\kappa_{1}\nabla u_{1}\cdot\nabla u_{1}$. The same steps hold for the second equation. Hence, summing the two equations results in the total energy for the two-membrane source problem
$J((u_{1},u_{2}))=\frac{1}{2}\int_{\Omega}\kappa_{1}\nabla u_{1}\cdot\nabla u_{1}-\int_{\Omega}f_{1}u_{1}+\frac{1}{2}\int_{\Omega}\kappa_{2}\nabla u_{2}\cdot\nabla u_{2}-\int_{\Omega}f_{2}u_{2}.$ (29)
Next, we analyze the continuous saddle point problem and find the corresponding Lagrange multiplier as we move towards the general Nitsche formulation (21). The constraint is given by $\beta((u_{1},u_{2}))=u_{2}-u_{1}+g\geq 0$. Hence, for the Lagrange multiplier $\lambda((u_{1},u_{2}))$, consider the Lagrangian
$\mathcal{L}((u_{1},u_{2}),\lambda)=J((u_{1},u_{2}))-\int_{\Omega}\lambda((u_{1},u_{2}))\beta((u_{1},u_{2})).$ (30)
Let $\kappa_{1}<\kappa_{2}$ such that the first membrane is less stiff. We can find the strong form of the Lagrange multiplier by setting the Gateaux derivative with respect to $v=(v_{1},0)$ to zero, and finally applying integration by parts:
$\mathcal{L}(u,\lambda;v)$ $=\bigg(\frac{1}{2}\int_{\Omega}\kappa_{1}\nabla(u_{1}+\epsilon v_{1})\cdot\nabla(u_{1}+\epsilon v_{1})-\int_{\Omega}f_{1}(u_{1}+\epsilon v_{1})$
$\quad\quad+\frac{1}{2}\int_{\Omega}\kappa_{2}\nabla u_{2}\cdot\nabla u_{2}-\int_{\Omega}f_{2}u_{2}-\int_{\Omega}\lambda(u_{2}-(u_{1}+\epsilon v_{1})+g)\bigg)\bigg|_{\epsilon=0}$
$=\bigg(\frac{1}{2}\int_{\Omega}\kappa_{1}(\nabla u_{1}\cdot\nabla u_{1}+2\epsilon\nabla u_{1}\cdot\nabla v_{1}+\epsilon^{2}\nabla v_{1}\cdot\nabla v_{1})-\int_{\Omega}f_{1}u_{1}+\epsilon f_{1}v_{1}$
$\quad\quad+\frac{1}{2}\int_{\Omega}\kappa_{2}\nabla u_{2}\cdot\nabla u_{2}-\int_{\Omega}f_{2}u_{2}-\int_{\Omega}\lambda(u_{2}-u_{1}-\epsilon v_{1}+g)\bigg)\bigg|_{\epsilon=0}$
$=\int_{\Omega}\kappa_{1}\nabla u_{1}\cdot\nabla v_{1}-f_{1}v_{1}+\lambda v_{1}$
$=\int_{\Omega}-\kappa_{1}\Delta u_{1}v_{1}-f_{1}v_{1}+\lambda v_{1}$
$=0$
$\implies\lambda$ $=\kappa_{1}\Delta u_{1}+f_{1}.$
We now have the constraint and an expression for the Lagrange multiplier, but must still ensure that the stabilized discrete problem has the correct units. As described in Section 4, $\gamma$ scales the discrete problem such that $\lambda\propto\frac{\beta}{\gamma}$. Because the discrete solution $u_{h}$ is differentiated twice in the definition $\lambda(u_{h})$ and $\beta(u_{h})$ includes the solution $u_{h}$ in its original units (i.e., in meters), $\gamma$ must here include a square of the mesh parameter. Further, as we derived $\lambda$ with respect to the less stiff side per [gustafsson2019error], the tension of the first membrane is included in $\lambda$ and correspondingly $\gamma$ must have its reciprocal. Hence, it must be that $\gamma(h,\kappa)=\frac{\alpha h^{2}}{\kappa_{1}}$ for the units to coincide.
To collect, the Nitsche terms are
$\lambda(u_{h})=\Delta_{h}u_{h,1}+f_{1},\quad\beta(u_{h})=u_{h,2}-u_{h,1}+g,\quad\gamma(h)=\frac{\alpha h^{2}}{\kappa_{1}},$
where $\Delta_{h}$ denotes the element-wise Laplacian operator. Substituting these to (21) produces the Nitsche problem: find $(u_{h,1},u_{h,2})\in V_{h}\times V_{h}$ which minimizes
$J((u_{h,1},u_{h,2}))$ $+\int_{\Omega}\frac{\alpha h^{2}}{2\kappa_{1}}\left(\Delta_{h}u_{h,1}+f_{1}-\frac{\kappa_{1}}{\alpha h^{2}}(u_{h,2}-u_{h,1}+g)\right)_{+}^{2}$ (31)
$-\int_{\Omega}\frac{\alpha h^{2}}{2\kappa_{1}}\left(\Delta_{h}u_{h,1}+f_{1}\right)^{2}.$
Remark 3
The Nitsche terms and formulation are derived similarly for other problems in Section 6.
The problem was solved numerically over $\Omega=(0,1)^{2}$ with the parameterization $g=0.05$, $f_{1}=1,f_{2}=0,\kappa_{1}=\kappa_{2}=1$ and $\alpha={1}\mathrm{e}{-2}$ using linear triangular elements. Note that the element-wise Laplacian for linear elements is equal to zero which simplifies the formulation. The numerical solution is displayed in Figure 1 and the approximated linear convergence rate in Figure 2.
Figure 2: Two membrane contact problem convergence rate follows the theoretical linear convergence in the $H^{1}$ norm for linear finite element basis.
As described earlier, we can revert to the penalty method by removing the terms depending on $\lambda((u_{1},u_{2}))$. However, the penalty method retains the optimal convergence rate only with linear elements, and to produce the same level of accuracy as Nitsche for quadratic or higher order elements we need to increase the power of $h$ in $\gamma$ [gustafsson2017finite]. This leads to higher condition numbers in the Newton iteration, see Figure 3 for an example in the quadratic case.
Figure 3: Condition numbers of the Jacobian in the Newton iterations (45) for the penalty and Nitsche variants of the two membrane contact problem with quadratic elements. The blue lines depict condition numbers of the penalty method and red lines of the Nitsche method, respectively. The markers distinguish problems of different sizes. Compared to the Nitsche method’s $\gamma=\alpha h^{2}$, the penalty method requires $\gamma=\alpha h^{3}$ to retain optimal convergence [gustafsson2017finite] and has thus consistently larger conditioning numbers with slower convergence.
6.2 Membrane against elastic solid
Let $\Omega_{1}$ describe the membrane
$\Omega_{1}=\{(x,y,0)\in\mathbb{R}^{3}:0<x,y<1\}$
and $\Omega_{2}$ describe the linear elastic cube, initially $g$ units apart from $\Omega_{1}$:
$\Omega_{2}=\{(x,y,z+g)\in\mathbb{R}^{3}:0<x,y,z<1\}.$
The energy of the source problem including membrane and linear elastic solid reads
$J((u_{1},\bm{u}_{2}))=J_{1}(u_{1})+J_{2}(\bm{u}_{2})$ (32)
where
$J_{1}(u_{1})=\int_{\Omega_{1}}\nabla u_{1}\cdot\nabla u_{1}-\int_{\Omega_{1}}fu_{1}$ (33)
and
$J_{2}(\bm{u}_{2})=\int_{\Omega_{2}}\bm{\sigma}(\bm{u}_{2}):\bm{\varepsilon}(\bm{u}_{2}).$ (34)
The functions $\bm{\sigma}$ and $\bm{\varepsilon}$ have been defined in equations (25) and (26), respectively.
Given the initial gap $g>0$, we set the physical non-interpenetration constraint as $u_{1}-\bm{u}_{2}\cdot\bm{n}\leq g$ (in $\Omega_{1}$) which results in $\beta((u_{1},\bm{u}_{2}))=\bm{u}_{2}\cdot\bm{n}-u_{1}+g\geq 0$. The Lagrange multiplier for the problem is $\lambda((u_{1},\bm{u}_{2}))=-\bm{\sigma}(\bm{u}_{2})\bm{n}_{2}\cdot\bm{n}_{2}$ and the scaling is $\gamma(h)=\alpha h$. Thus, the Nitsche method minimizes the constrained energy
$J(u_{1,h},\bm{u}_{2,h})$ $+\int_{\Omega_{1}}\frac{\alpha h}{2}\left(-\bm{\sigma}(\bm{u}_{2})\bm{n}_{2}\cdot\bm{n}_{2}-\frac{1}{\alpha h}(\bm{u}_{2}\cdot\bm{n}-u_{1}+g)\right)_{+}^{2}$ (35)
$-\int_{\Omega_{1}}\frac{\alpha h}{2}(-\bm{\sigma}(\bm{u}_{2})\bm{n}_{2}\cdot\bm{n}_{2})^{2}.$
The problem was solved numerically using a gap of $g=0.1$, $\mu_{L}=\lambda_{L}=1$, $f=2$ and $\alpha={1}\mathrm{e}{-2}$ using linear hexahedral elements for the solid and linear quadrilateral elements for the membrane. The numerical solution is displayed in Figure 4 and the approximated linear convergence rate in Figure 5. The first principal stress for the solid is presented in Figure 6.
Figure 4: Numerical solution for the membrane against elastic solid problem using the Nitsche method.
Figure 5: Convergence rate in the $H^{1}$ norm for the membrane against elastic solid problem.
Figure 6: The first principal stress for the elastic solid in membrane against elastic solid problem.
6.3 Plate against plate
The plate against plate problem follows very closely the two membrance contact problem with the difference of higher order derivatives in the strain energy. The energy of the source problem is
$J((u_{1},u_{2}))=\frac{1}{2}\int_{\Omega}\sum_{i,j=1}^{2}\left(\frac{\partial u_{1}}{\partial x_{i}x_{j}}\right)^{2}-\int_{\Omega}f_{1}u_{1}+\frac{1}{2}\int_{\Omega}\sum_{i,j=1}^{2}\left(\frac{\partial u_{2}}{\partial x_{i}x_{j}}\right)^{2}-\int_{\Omega}f_{2}u_{2}.$ (36)
We assume again a gap $g>0$ such that $u_{1}-u_{2}\leq g$ in $\Omega$ for the constraint $\beta{(u_{1},u_{2})}=u_{2}-u_{1}+g\geq 0$. Then, we have $\lambda((u_{1},u_{2}))=-\Delta^{2}u_{1}-f_{1}$ and $\gamma(h)=\alpha h^{4}$.
Hence, the Nitsche problem minimizes the energy
$J((u_{h,1},u_{h,2}))$ $+\int_{\Omega}\frac{\alpha h^{4}}{2}\left(-\Delta_{h}^{2}u_{h,1}-f_{1}-\frac{1}{\alpha h^{4}}(u_{h,2}-u_{h,1}+g)\right)_{+}^{2}$ (37)
$-\int_{\Omega}\frac{\alpha h^{4}}{2}\left(-\Delta_{h}^{2}u_{h,1}-f_{1}\right)^{2},$
where $\Delta_{h}^{2}$ denotes the element-wise biharmonic operator.
The problem was solved numerically over $\Omega=(0,1)^{2}$, a gap of $g=0.05$, $f_{1}=100$, $f_{2}=0$ and $\alpha={1}\mathrm{e}{-2}$ using triangular Morley elements [beirao2010posteriori]. Notice that the Morley element is quadratic and hence the element-wise biharmonic operator evaluates to zero. The numerical solution is displayed in Figure 7 and the approximated linear convergence rate in Figure 8.
Figure 7: Numerical solution for two plates in contact using the Nitsche method. The nonconforming Morley element results in small discontinueties over element boundaries.
Figure 8: Two plate contact problem convergence rate follows the theoretical linear convergence in the $H^{2}$ norm with the Morley elements [beirao2010posteriori].
6.4 Kirchhoff plate with inequality boundary condition
Let $\Omega=(0,1)^{2}$. For the plate corners problem [nazarov2012hinged], the energy with normalized units reads
$J(u)=\int_{\Omega}\sum_{i,j=1}^{2}\left(\frac{\partial^{2}u}{\partial x_{i}\partial x_{j}}\right)^{2}-\int_{\Omega}fu.$ (38)
We consider the problem where $\Gamma=\partial\Omega$ and $\beta(u)=u\geq 0$ on $\Gamma$. Then, the scaling parameter is $\gamma(h)=\alpha h^{3}$ and the Lagrange multiplier is $\lambda(u)=K(u):=Q_{n}+\frac{\partial M_{ns}}{\partial s}$, where M=HuM=-Hu, $Q=\text{div}\>M$, $n$ is the normal vector and $s$ the tangent vector on $\Gamma$ and $K(u)$ is called the Kirchhoff shear force [gustafsson2021nitsche].
The Nitsche functional is thus
$J(u_{h})+\int_{\Gamma}\frac{\alpha h^{3}}{2}\Big(K_{h}(u)-\frac{1}{\alpha h^{3}}u_{h}\Big)_{+}^{2}-\int_{\Gamma}\frac{\alpha h^{3}}{2}K_{h}(u)^{2},$ (39)
where $K_{h}$ is the element-wise evaluation of $K$.
The problem was solved numerically with f=1f=-1 and $\alpha=1/4$ using quadrilateral Bogner–Fox–Schmit elements [bogner1965generation]. The numerical solution is displayed in Figure 9 and the approximated quadratic convergence rate in Figure 10.
Figure 9: The corners of the plate are displaced upwards when downwards point load is applied at the middle of the plate, approximate solution using Nitsche’s method.
Figure 10: The convergence rate of the Kirchhoff plate with inequality boundary condition follows the theoretical quadratic convergence.
7 Conclusions
While the Nitsche method is commonly viewed as a consistency correction to the penalty method, we demonstrate how new methods can be obtained starting from a general minimization form. This allows for developing novel Nitsche methods for problems with both equality and inequality constraints in a simple manner. We derived and implemented several new problems using the approach. Our numerical experiments suggest that the methods converge optimally while the more careful numerical analysis remains a topic of future work.
Acknowledgements
We wish to thank Rolf Stenberg for valuable comments on the manuscript.
Appendix A Newton’s method
Consider a problem-dependent Sobolev space $U$ and its corresponding subspace $V$ of functions with zero boundary values so that $V\subset U$ and $u+v\in U\>\>\>\forall v\in V$. We want to find a solution to the energy minimization problem
$\min_{u\in U}J(u)$ (40)
where $J:U\to\mathbb{R}$ represents the energy of the system. Let $v\in V$ denote an arbitrary perturbation to $u$. A solution $u\in U$ to (40) satisfies
$J^{\prime}(u;v)=0\quad\forall v\in V,$ (41)
where $J^{\prime}:U\times V\to\mathbb{R}$ denotes the directional (Gateaux) derivative defined by
$J^{\prime}(u;v)=\frac{\partial J(u+\epsilon v)}{\partial\epsilon}\bigg|_{\epsilon=0}.$
The directional derivative is linear in $v$, but, in general, nonlinear in $u$.
Let $f=J^{\prime}$. We assume that the problem
$f(u;v)=0\quad\forall v\in V$
is nonlinear in $u$ and, hence, the standard Ritz–Galerkin method cannot be directly applied to obtain a linear system of equations. Consequently, we approach the problem as a root-finding problem and linearize $f$ to use the Newton’s method. More precisely, let $u_{0}\in U$ be an initial value such that the solution is written as $u=u_{0}+w$ for some $w\in V$. The perturbation $w$ is unknown, but can be approximated by discarding the higher order terms of the Taylor series of $f$:
$0=f(u;v)=f(u_{0}+w;v)\approx f(u_{0};v)+f^{\prime}(u_{0};v,w),$
where $f^{\prime}:U\times V\times V\to\mathbb{R}$ is the directional derivative of $f$.
A single step of the Newton’s method then reads: find $w\in V$ such that
$f^{\prime}(u_{0};v,w)=-f(u_{0};v)\quad\forall v\in V.$ (42)
Note that if $u_{0}\in U$ is fixed, then the problem is linear with respect to $w\in V$ and the standard Ritz–Galerkin method can be applied iteratively.
For the numerical scheme, we use the finite element method. Let $\mathcal{T}_{h}$ be a mesh of elements with maximum edge length $h>0$. The finite dimensional subspaces $U_{h}\subset U$ and $V_{h}\subset V$ contain element-wise polynomial functions of a given degree. The discrete problem then reads: given $u_{0,h}\in U_{h}$, find $w_{h}\in V_{h}$ such that
$f^{\prime}(u_{0,h};v_{h},w_{h})=-f(u_{0,h};v_{h})\quad\forall v_{h}\in V_{h}.$ (43)
The directional derivative of $f$ is computable as $u_{0,h}$ is given.
The equation (43) can be posed as a linear system of equations. Let $\{\phi_{i}\}_{i=1}^{n}$, $n=\dim(V_{h})$, be a finite element basis that span $V_{h}$. Then
$w_{h}=\sum_{j=1}^{n}\beta_{j}\phi_{j},$ (44)
where $\beta_{j}\in\mathbb{R}$ are the unknown degrees-of-freedom. Choosing $v_{h}=\phi_{i}$, we have
$\sum_{j=1}^{n}\beta_{j}f^{\prime}(u_{0,h};\phi_{i},\phi_{j})=-f(u_{0,h};\phi_{i}),$
and collecting all $i=1,\dots,n$ results in a linear system
$\bm{J}\bm{\beta}=\bm{f}$ (45)
where
$\bm{J}_{ij}=f^{\prime}(u_{0,h};\phi_{i},\phi_{j}),\quad\bm{f}_{i}=-f(u_{0,h};\phi_{i}),$
or, if the problem is given directly in the minimization form,
$\bm{J}_{ij}=J^{\prime\prime}(u_{0,h};\phi_{i},\phi_{j}),\quad\bm{f}_{i}=-J^{\prime}(u_{0,h};\phi_{i}).$
The direction (44) is then found from the solution $\bm{\beta}$.
To conclude, starting from an initial value $u_{0,h}\in U_{h}$, we find an approximation $u_{h}\in U_{h}$ such that $J^{\prime}(u_{h};v_{h})=0$ for all $v_{h}\in V_{h}$ through the iterative scheme $u_{k+1,h}=u_{k,h}+\alpha_{k}w_{k,h}$, where $0<\alpha_{k}\leq 1$ is a step size parameter. Under standard assumptions the method converges quadratically.